# Convergence of the open WASEP stationary measure without Liggett’s condition

### v1

@@ 3.1. Convergence of One-point Distributions
$-\log{\left(4\pi N\sqrt{\frac{r}{N}}\right)}+\mathcal{A}_{N}(\sqrt{m},\sqrt{r})=$
$\log{\left(\frac{\bigg{|}\Gamma\left(v+\frac{s}{2}+i\frac{\sqrt{m}}{2}\right)\Gamma\left(\frac{t-s}{2}+i\frac{\sqrt{r}+\sqrt{m}}{2}\right)\Gamma\left(\frac{t-s}{2}+i\frac{\sqrt{r}-\sqrt{m}}{2}\right)\bigg{|}^{2}}{8\pi\sqrt{r}\Gamma(t-s)\bigg{|}\Gamma\left(v+\frac{t+i\sqrt{r}}{2}\right)\Gamma\left(\frac{i\sqrt{m}}{2}\right)\bigg{|}^{2}}\right)}$
We also simplify the expression for the ratio of the one-point distributions which arises in (3.9).
$\log{\left(\frac{\sqrt{1-\frac{r}{4N}}\hat{\pi}_{t;v,w,u,r}^{(N),c}(r)}{\sqrt{1-\frac{m}{4N}}\hat{\pi}_{s;v,w,u,r}^{(N),c}(m)}\right)}=\log{\left(\frac{\mathfrak{p}_{t}^{c}(r)}{\mathfrak{p}_{s}^{c}(m)}\right)}+\textsc{Error}_{t}^{(N),c}(r)-\textsc{Error}_{s}^{(N),c}(m)$
@@
$=\log{\left(\mathfrak{p}_{s,t}^{c,c}(m,r)\right)}+E_{N}^{\mathcal{A}}(m,r)+E_{N}^{\mathcal{E}}(m,r)+\textsc{Error}_{t}^{(N),c}(r)-\textsc{Error}_{s}^{(N),c}(m)$
@@
For all $\eta>1$ there exists $N_{0}\in\mathbb{Z}_{\geq 1}$ and $C,\chi>0$ such that for all $N>N_{0}$ and $m,r\in\hat{S}^{(N),c}=[0,4N]$,
$|E_{N}^{\mathcal{A}}(m,r)|,|E_{N}^{\mathcal{E}}(m,r)|\leq CN^{-\chi}(1+\sqrt{m})^{\eta}+CN^{-\chi}(1+\sqrt{r})^{\eta}$
@@
We rewrite $\hat{\pi}_{s,t;v,w,u,r}^{(N),d,c}\left(\hat{y}_{j}^{(N),\chi}(s),r\right)$ through the same time reversal transformation:
(3.11) $\hat{\pi}_{s,t;v,w,u,r}^{(N),d,c}(\hat{y}_{j}^{(N),\chi}(s),r)=\pi_{q^{t},q^{s}}^{(N),c,d}\left(1-\frac{r}{2N},y_{j}^{(N),\chi}(s)\right)\cdot\frac{\hat{\pi}_{t;v,w,u,r}^{(N),c}(r)}{\hat{\pi}_{s}^{(N),d}(\hat{y}_{j}^{(N),\chi}(s))}$
@@
$\pi_{q^{t},q^{s}}^{(N)}\left(1-\frac{r}{2N},V\right)=AW(V;q^{v+\frac{s}{2}},-q^{w+\frac{s}{2}},q^{\frac{t-s}{2}+i\frac{\sqrt{N}\theta_{r}}{2}},q^{\frac{t-s}{2}-i\frac{\sqrt{N}\theta_{r}}{2}})$
@@
(3.12) $\pi_{q^{t},q^{s}}^{(N),c,d}\left(1-\frac{r}{2N},y_{0}^{(N),v}(s)\right)$ $=\frac{(q^{-2v-s},-q^{w+\frac{t}{2}+i\frac{\sqrt{N}\theta_{r}}{2}},-q^{w+\frac{t}{2}-i\frac{\sqrt{N}\theta_{r}}{2}},q^{t-s})_{\infty}}{(-q^{w-v},q^{-v+\frac{t}{2}-s+i\frac{\sqrt{N}\theta_{r}}{2}},q^{-v+\frac{t}{2}-s-i\frac{\sqrt{N}\theta_{r}}{2}},-q^{v+t+w})_{\infty}},$
(3.13) $\pi_{q^{t},q^{s}}^{(N),c,d}\left(1-\frac{r}{2N},y_{j}^{(N),v}(s)\right)$ $=\pi_{q^{t},q^{s}}^{(N),c,d}\left(1-\frac{r}{2N},y_{0}^{(N),v}(s)\right)\times M_{j},$
(3.14) $M_{j}$ $=\frac{(q^{2v+s},-q^{v+w+s},q^{v+\frac{t}{2}+i\frac{\sqrt{N}\theta_{r}}{2}},-q^{v+\frac{t}{2}-i\frac{\sqrt{N}\theta_{r}}{2}})_{j}(1-q^{2v+2j+s})}{(q,-q^{1+v-w},q^{1+v+s-\frac{t}{2}-i\frac{\sqrt{N}\theta_{r}}{2}},q^{1+v+s-\frac{t}{2}+i\frac{\sqrt{N}\theta_{r}}{2}})_{j}(1-q^{2v+s})(-q^{v+w+t-1})^{j}}.$
@@
$\log{\left(\pi_{q^{t},q^{s}}^{(N),c,d}\left(1-\frac{r}{2N},y_{0}^{(N),v}(s)\right)\right)}=\mathcal{A}_{N}(\sqrt{r})+E_{N}^{\mathcal{A}}(r)+E_{N}^{\mathcal{E}}(r)$
@@
(3.15) $\mathcal{A}_{N}(\sqrt{r})$ $:=\mathcal{A}_{N}^{+}\left[-2v-s,t-s\right]+\mathcal{A}_{N}^{-}\left[b+\frac{t}{2}+i\frac{\sqrt{r}}{2},b+\frac{t}{2}-i\frac{\sqrt{r}}{2}\right]$
$-\mathcal{A}_{N}^{+}\left[-v+\frac{t}{2}-s+i\frac{\sqrt{r}}{2},-v+\frac{t}{2}-s-i\frac{\sqrt{r}}{2}\right]-\mathcal{A}_{N}^{-}\left[b-v,v+t+b\right]$
$=\log{\left(\frac{\bigg{|}\Gamma\left(\frac{t}{2}-s-v+\frac{i\sqrt{r}}{2}\right)\bigg{|}^{2}}{\Gamma(-2v-s)\Gamma(t-s)}\right)}$
@@
We also need to address the ratio of $\hat{\pi}_{t}^{(N),c}(r)$ and $\hat{\pi}_{s}^{(N),d}(y_{j}^{(N),v}(s))$ which arose in (3.11).
$\log{\left(\frac{\sqrt{1-\frac{r}{4N}}\hat{\pi}_{t;v,w,u,r}^{(N),c}(r)}{\hat{\pi}_{s}^{(N),d}(\hat{y}_{j}^{(N),v}(s))}\right)}=\log{\left(\frac{\mathfrak{p}_{t}^{c}(y)}{\mathfrak{p}_{s}^{d}(x_{j}^{v})}\right)}+\textsc{Error}_{t}^{(N),c}(r)-\textsc{Error}_{s}^{(N),d}(x_{0}^{v}(s))$
@@
(3.16) $\log{\left(\sqrt{1-\frac{r}{4N}}\hat{\pi}_{s,t;v,w,u,r}^{(N),d,c}(\hat{y}_{0}^{(N),v}(s),r)\right)}$ $=\log{\left(\frac{|\Gamma(\frac{t}{2}-s-v+\frac{i\sqrt{r}}{2})|^{2}\mathfrak{p}_{t}^{c}(r)}{\Gamma(-2v-s)\Gamma(t-s)\mathfrak{p}_{s}^{d}(x_{j}^{v})}\right)}+E_{N}^{\mathcal{A}}(r)$
@@
$+\textsc{Error}_{t}^{(N),c}(r)-\textsc{Error}_{s}^{(N),d}(x_{0}^{v}(s))$
@@
For $E_{N}^{\mathcal{A}}(r)$ and $E_{N}^{\mathcal{E}}(r)$ as defined by (3.15), for all $\eta>1$ there exists $N_{0}\in\mathbb{Z}_{\geq 1}$ and $C,\chi>0$ such that for all $N>N_{0}$ and $r\in\hat{S}^{(N),c}=[0,4N]$,
$|E_{N}^{\mathcal{A}}(r)|,|E_{N}^{\mathcal{E}}(r)|\leq CN^{-\chi}(1+\sqrt{r})^{\eta}$
@@
(3.17) $M_{j}=\frac{[2v+s,v+\frac{t}{2}+i\frac{\sqrt{r}}{2},v+\frac{t}{2}-i\frac{\sqrt{r}}{2}]_{j}(2v+2j+s)}{[1,1+v+s-\frac{t}{2}+i\frac{\sqrt{r}}{2},1+v+s-\frac{t}{2}-i\frac{\sqrt{r}}{2}]_{j}(2v+s)}(-1)^{j}e^{E_{N}^{M_{j}}(r)}$
Where, for all $\eta>1$ there exists $N_{0}\in\mathbb{Z}_{\geq 1}$ and $C,\chi>0$ such that for all $N>N_{0}$ and $r\in\hat{S}_{t}^{(N),c}=[0,4N]$,
$|E_{N}^{M_{j}}(r)|\leq CN^{-\chi}(1+\sqrt{r})^{\eta}$
@@
$\log{\left(\sqrt{1-\frac{r}{4N}}\hat{\pi}_{s,t;v,w,u,r}^{(N),d,c}(\hat{y}_{j}^{(N),v}(s),r)\right)}=\log{\left(\mathfrak{p}_{s,t}^{d,c}(x_{j}^{v},r)\right)}+\overline{\textsc{Error}}_{s,t}^{(N),d,c}(x_{j}^{v}(s),r)$
With $\overline{\textsc{Error}}_{s,t}^{(N),d,c}(x_{j}^{v}(s),r):=\textsc{Error}_{s,t}^{(N),d,c}(x_{j}^{v}(s),r)+E_{N}^{M_{j}}$. By the above calculations, for all $\eta>1$, there exist $N_{0}\in\mathbb{Z}_{\geq 1}$ and $C,\chi>0$ such that for all $N>N_{0}$, $r\in\hat{S}^{(N),c}_{t}$,
$\overline{\textsc{Error}}_{s,t}^{(N),d,c}(x_{j}^{v}(s),r)\leq CN^{-\chi}(1+\sqrt{r})^{\eta}$
@@
$\pi_{q^{t},q^{s}}^{(N)}\left(1-\frac{r}{2N},V\right)=AW(V;q^{v+\frac{s}{2}},-q^{w+\frac{s}{2}},q^{\frac{t-s}{2}+i\frac{\sqrt{N}\theta_{r}}{2}},q^{\frac{t-s}{2}-i\frac{\sqrt{N}\theta_{r}}{2}})$
@@
$\hat{\pi}_{s,t}^{(N),d,d}(x,y)=\pi_{q^{t},q^{s}}^{(N),d,d}\left(1-\frac{y}{2N},1-\frac{x}{2N}\right)\cdot\frac{\hat{\pi}_{t}^{(N),d}(y)}{\hat{\pi}_{s}^{(N),d}(x)}$
@@
$\pi_{q^{t},q^{s}}^{(N)}\left(y_{j}^{(N),v}(t),V\right)$ $=AW\left(V;q^{v+\frac{s}{2}},-q^{w+\frac{s}{2}},q^{\frac{t}{2}+v+j},q^{\frac{t}{2}-s-v-j}\right)$
@@
$\pi_{q^{t},q^{s}}^{(N)}\left(y_{j}^{(N),v}(t),y_{0}^{(N),v}(s)\right)$ $=\frac{(q^{-2v-s},-q^{w+t+j+v},-q^{w-j-v},q^{t-s})_{\infty}}{(-q^{w-v},q^{t-s+j},q^{-2v-s-j},-q^{v+w+t})_{\infty}}$
$\pi_{q^{t},q^{s}}^{(N)}\left(y_{j}^{(N),v},y_{k}^{(N),v}\right)$ $=\pi_{q^{t},q^{s}}^{(N)}\left(y_{j}^{(N),v},y_{0}^{(N),v}\right)\times M_{k}$
$M_{k}$ $=\frac{(q^{2v+s},-q^{v+w+s},q^{2v+t+j},q^{-j})_{k}(1-q^{2v+2k+s})}{(q,-q^{1+v-w},q^{1+s-t-j},q^{1+2v+s+j})_{k}(1-q^{2v+s})(-q^{v+w+t-1})^{k}}$
@@
$\mathcal{A}_{N}(j)$ $:=\mathcal{A}^{+}_{N}[-2v-s,t-s]+\mathcal{A}^{-}[w+t+j+v,w-j-v]$
$-\mathcal{A}^{+}[t-s+j,-2v-s-j]-\mathcal{A}^{-}[w-v,v+w+t]$
@@
$\pi_{q^{t},q^{s}}^{(N),d,d}\left(y_{j}^{(N),v}(t),y_{0}^{(N),v}(s)\right)=\frac{\Gamma(t-s+j)\Gamma(-2v-s-j)}{\Gamma(-2v-s)\Gamma(t-s)}e^{\textsc{Error}_{t,s}^{(N),d,d}(y_{j}^{(N),v}(t),y_{0}^{(N),v}(s))}$
@@
$M_{k}=\frac{[2v+s,2v+t+j,-j]_{k}(2v+2k+s)}{[1,1+s-t-j,1+2v+s+j]_{k}(2v+s)}(-1)^{k}e^{\overline{\textsc{Error}}_{t,s}^{(N),d,d}(y_{j}^{(N),v}(t),y_{k}^{(N),v}(s))}$
@@
$|\overline{\textsc{Error}}_{t,s}^{(N),d,d}(y_{j}^{(N),v}(t),y_{k}^{(N),v}(s))|\leq CN^{-\chi}$
This finishes the proof of Lemma 3.4, part (1). ∎
We can also deal with the limit of the ration of $\hat{\pi}_{t}^{(N),d}(y)$ and $\hat{\pi}_{s}^{(N),d}(x)$ using our one-point computations.
$\lim_{N\to\infty}\frac{\hat{\pi}_{t}^{(N),d}(\hat{y}_{j}^{(N),v})}{\hat{\pi}_{s}^{(N),d}(\hat{y}_{j}^{(N),v})}$ $=\lim_{N\to\infty}\frac{N^{u+v}\hat{\pi}_{t}^{(N),d}(\hat{y}_{j}^{(N),v}(t))}{N^{u+v}\hat{\pi}_{s}^{(N),d}(\hat{y}_{k}^{(N),v}(s))}=\frac{\mathfrak{p}_{t}^{d}(x_{j}^{v}(t))}{\mathfrak{p}_{s}^{d}(x_{j}^{v}(s))}$
@@
$\lim_{N\to\infty}\hat{\pi}_{s,t}^{(N),d,d}(\hat{y}_{j}^{(N),v},\hat{y}_{k}^{(N),v})=\mathfrak{p}_{s,t}(x_{k}^{v}(s),x_{j}^{v}(t))$
@@
$\hat{\pi}_{s,t}^{(N),d,d}\left(\hat{y}_{j}^{(N),u}(s),\hat{y}_{k}^{(N),v}(t)\right)$ $=\pi_{q^{t},q^{s}}^{(N),d,d}\left(y_{k}^{(N),v}(t),y_{j}^{(N),u}(s)\right)\frac{\pi_{q^{t}}^{(N),d}\left(y_{k}^{(N),v}(t)\right)}{\pi_{q^{s}}^{(N),d}\left(y_{j}^{(N),u}(s)\right)}$
$\pi_{q^{t},q^{s}}^{(N),d,c}\left(y_{k}^{(N),v}(t),V\right)$ $=AW\left(V;q^{v+\frac{s}{2}},-q^{w+\frac{s}{2}},q^{\frac{t-s}{2}+\left(k+v+\frac{t}{2}\right)},q^{\frac{t-s}{2}-\left(k+v+\frac{t}{2}\right)}\right)$

### v2

@@ 3.1. Convergence of Marginal Distributions
$-\log{\left(4\pi N\sqrt{\frac{r}{N}}\right)}+\mathcal{A}_{N}(\sqrt{m},\sqrt{r})$
$=\log{\left(\frac{\bigg{|}\Gamma\left(v+\frac{s}{2}+i\frac{\sqrt{m}}{2}\right)\Gamma\left(\frac{t-s}{2}+i\frac{\sqrt{r}+\sqrt{m}}{2}\right)\Gamma\left(\frac{t-s}{2}+i\frac{\sqrt{r}-\sqrt{m}}{2}\right)\bigg{|}^{2}}{8\pi\sqrt{r}\Gamma(t-s)\bigg{|}\Gamma\left(v+\frac{t+i\sqrt{r}}{2}\right)\Gamma\left(\frac{i\sqrt{m}}{2}\right)\bigg{|}^{2}}\right)}.$
We also simplify the expression for the ratio of the marginal distributions which arises in (3.9).
$\log{\left(\frac{\sqrt{1-\frac{r}{4N}}\widehat{\pi}_{t;v,w,u,r}^{(N),c}(r)}{\sqrt{1-\frac{m}{4N}}\widehat{\pi}_{s;v,w,u,r}^{(N),c}(m)}\right)}=\log{\left(\frac{\mathfrak{p}_{t}^{c}(r)}{\mathfrak{p}_{s}^{c}(m)}\right)}+\textsc{Error}_{t}^{(N),c}(r)-\textsc{Error}_{s}^{(N),c}(m),$
@@
$=\log{\left(\mathfrak{p}_{s,t}^{c,c}(m,r)\right)}+E_{N}^{\mathcal{A}}(m,r)+E_{N}^{\mathcal{E}}(m,r)+\textsc{Error}_{t}^{(N),c}(r)-\textsc{Error}_{s}^{(N),c}(m).$
@@
For all $\eta>1$ there exists $N_{0}\in\mathbb{Z}_{\geq 1}$ and $C,\chi>0$ such that for all $N>N_{0}$ and $m,r\in\widehat{S}^{(N),c}=[0,4N]$,
$|E_{N}^{\mathcal{A}}(m,r)|,|E_{N}^{\mathcal{E}}(m,r)|\leq CN^{-\chi}(1+\sqrt{m})^{\eta}+CN^{-\chi}(1+\sqrt{r})^{\eta}.$
@@
We rewrite $\widehat{\pi}_{s,t;v,w,u,r}^{(N),d,c}\left(\widehat{y}_{j}^{(N),\chi}(s),r\right)$ through the same time reversal transformation:
$\widehat{\pi}_{s,t;v,w,u,r}^{(N),d,c}(\widehat{y}_{j}^{(N),\chi}(s),r)=\pi_{q^{t},q^{s}}^{(N),c,d}\left(1-\frac{r}{2N},y_{j}^{(N),\chi}(s)\right)\cdot\frac{\widehat{\pi}_{t;v,w,u,r}^{(N),c}(r)}{\widehat{\pi}_{s}^{(N),d}(\widehat{y}_{j}^{(N),\chi}(s))}.$ (3.11)
@@
$\pi_{q^{t},q^{s}}^{(N)}\left(1-\frac{r}{2N},V\right)=AW(V;q^{v+\frac{s}{2}},-q^{w+\frac{s}{2}},q^{\frac{t-s}{2}+i\frac{\sqrt{N}\theta_{r}}{2}},q^{\frac{t-s}{2}-i\frac{\sqrt{N}\theta_{r}}{2}}).$
@@
$\pi_{q^{t},q^{s}}^{(N),c,d}\left(1-\frac{r}{2N},y_{0}^{(N),v}(s)\right)$ $=\frac{(q^{-2v-s},-q^{w+\frac{t}{2}+i\frac{\sqrt{N}\theta_{r}}{2}},-q^{w+\frac{t}{2}-i\frac{\sqrt{N}\theta_{r}}{2}},q^{t-s})_{\infty}}{(-q^{w-v},q^{-v+\frac{t}{2}-s+i\frac{\sqrt{N}\theta_{r}}{2}},q^{-v+\frac{t}{2}-s-i\frac{\sqrt{N}\theta_{r}}{2}},-q^{v+t+w})_{\infty}},$ (3.12)
$\pi_{q^{t},q^{s}}^{(N),c,d}\left(1-\frac{r}{2N},y_{j}^{(N),v}(s)\right)$ $=\pi_{q^{t},q^{s}}^{(N),c,d}\left(1-\frac{r}{2N},y_{0}^{(N),v}(s)\right)\times M_{j},$ (3.13)
$M_{j}$ $=\frac{(q^{2v+s},-q^{v+w+s},q^{v+\frac{t}{2}+i\frac{\sqrt{N}\theta_{r}}{2}},-q^{v+\frac{t}{2}-i\frac{\sqrt{N}\theta_{r}}{2}})_{j}(1-q^{2v+2j+s})}{(q,-q^{1+v-w},q^{1+v+s-\frac{t}{2}-i\frac{\sqrt{N}\theta_{r}}{2}},q^{1+v+s-\frac{t}{2}+i\frac{\sqrt{N}\theta_{r}}{2}})_{j}(1-q^{2v+s})(-q^{v+w+t-1})^{j}}.$ (3.14)
@@
$\log{\left(\pi_{q^{t},q^{s}}^{(N),c,d}\left(1-\frac{r}{2N},y_{0}^{(N),v}(s)\right)\right)}=\mathcal{A}_{N}(\sqrt{r})+E_{N}^{\mathcal{A}}(r)+E_{N}^{\mathcal{E}}(r).$
@@
$\mathcal{A}_{N}(\sqrt{r})$ $:=\mathcal{A}_{N}^{+}\left(-2v-s,t-s\right)+\mathcal{A}_{N}^{-}\left(b+\frac{t}{2}+i\frac{\sqrt{r}}{2},b+\frac{t}{2}-i\frac{\sqrt{r}}{2}\right)$ (3.15)
$-\mathcal{A}_{N}^{+}\left(-v+\frac{t}{2}-s+i\frac{\sqrt{r}}{2},-v+\frac{t}{2}-s-i\frac{\sqrt{r}}{2}\right)-\mathcal{A}_{N}^{-}\left(b-v,v+t+b\right)$
$=\log{\left(\frac{\bigg{|}\Gamma\left(\frac{t}{2}-s-v+\frac{i\sqrt{r}}{2}\right)\bigg{|}^{2}}{\Gamma(-2v-s)\Gamma(t-s)}\right)},$
@@
We also need to address the ratio of $\widehat{\pi}_{t}^{(N),c}(r)$ and $\widehat{\pi}_{s}^{(N),d}(y_{j}^{(N),v}(s))$ which arose in (3.11).
$\log{\left(\frac{\sqrt{1-\frac{r}{4N}}\widehat{\pi}_{t;v,w,u,r}^{(N),c}(r)}{\widehat{\pi}_{s}^{(N),d}(\widehat{y}_{j}^{(N),v}(s))}\right)}=\log{\left(\frac{\mathfrak{p}_{t}^{c}(y)}{\mathfrak{p}_{s}^{d}(x_{j}^{v})}\right)}+\textsc{Error}_{t}^{(N),c}(r)-\textsc{Error}_{s}^{(N),d}(x_{0}^{v}(s)).$
@@
$\log{\left(\sqrt{1-\frac{r}{4N}}\widehat{\pi}_{s,t;v,w,u,r}^{(N),d,c}(\widehat{y}_{0}^{(N),v}(s),r)\right)}$ $=\log{\left(\frac{|\Gamma(\frac{t}{2}-s-v+\frac{i\sqrt{r}}{2})|^{2}\mathfrak{p}_{t}^{c}(r)}{\Gamma(-2v-s)\Gamma(t-s)\mathfrak{p}_{s}^{d}(x_{j}^{v})}\right)}+E_{N}^{\mathcal{A}}(r)$ (3.16)
@@
$+\textsc{Error}_{t}^{(N),c}(r)-\textsc{Error}_{s}^{(N),d}(x_{0}^{v}(s)).$
@@
For $E_{N}^{\mathcal{A}}(r)$ and $E_{N}^{\mathcal{E}}(r)$ as defined by (3.15), for all $\eta>1$ there exists $N_{0}\in\mathbb{Z}_{\geq 1}$ and $C,\chi>0$ such that for all $N>N_{0}$ and $r\in\widehat{S}^{(N),c}=[0,4N]$,
$|E_{N}^{\mathcal{A}}(r)|,|E_{N}^{\mathcal{E}}(r)|\leq CN^{-\chi}(1+\sqrt{r})^{\eta}.$
@@
$M_{j}=\frac{[2v+s,v+\frac{t}{2}+i\frac{\sqrt{r}}{2},v+\frac{t}{2}-i\frac{\sqrt{r}}{2}]_{j}(2v+2j+s)}{[1,1+v+s-\frac{t}{2}+i\frac{\sqrt{r}}{2},1+v+s-\frac{t}{2}-i\frac{\sqrt{r}}{2}]_{j}(2v+s)}(-1)^{j}e^{E_{N}^{M_{j}}(r)}.$ (3.17)
Where, for all $\eta>1$ there exists $N_{0}\in\mathbb{Z}_{\geq 1}$ and $C,\chi>0$ such that for all $N>N_{0}$ and $r\in\widehat{S}_{t}^{(N),c}=[0,4N]$,
$|E_{N}^{M_{j}}(r)|\leq CN^{-\chi}(1+\sqrt{r})^{\eta}.$
@@
$\log{\left(\sqrt{1-\frac{r}{4N}}\widehat{\pi}_{s,t;v,w,u,r}^{(N),d,c}(\widehat{y}_{j}^{(N),v}(s),r)\right)}=\log{\left(\mathfrak{p}_{s,t}^{d,c}(x_{j}^{v},r)\right)}+\overline{\textsc{Error}}_{s,t}^{(N),d,c}(x_{j}^{v}(s),r),$
with $\overline{\textsc{Error}}_{s,t}^{(N),d,c}(x_{j}^{v}(s),r):=\textsc{Error}_{s,t}^{(N),d,c}(x_{j}^{v}(s),r)+E_{N}^{M_{j}}$. By the above calculations, for all $\eta>1$, there exist $N_{0}\in\mathbb{Z}_{\geq 1}$ and $C,\chi>0$ such that for all $N>N_{0}$, $r\in\widehat{S}^{(N),c}_{t}$,
$\overline{\textsc{Error}}_{s,t}^{(N),d,c}(x_{j}^{v}(s),r)\leq CN^{-\chi}(1+\sqrt{r})^{\eta}.$
@@
$\pi_{q^{t},q^{s}}^{(N)}\left(1-\frac{r}{2N},V\right)=AW(V;q^{v+\frac{s}{2}},-q^{w+\frac{s}{2}},q^{\frac{t-s}{2}+i\frac{\sqrt{N}\theta_{r}}{2}},q^{\frac{t-s}{2}-i\frac{\sqrt{N}\theta_{r}}{2}}).$
@@
$\widehat{\pi}_{s,t}^{(N),d,d}(x,y)=\pi_{q^{t},q^{s}}^{(N),d,d}\left(1-\frac{y}{2N},1-\frac{x}{2N}\right)\cdot\frac{\widehat{\pi}_{t}^{(N),d}(y)}{\widehat{\pi}_{s}^{(N),d}(x)}.$
@@
$\pi_{q^{t},q^{s}}^{(N)}\left(y_{j}^{(N),v}(t),V\right)$ $=AW\left(V;q^{v+\frac{s}{2}},-q^{w+\frac{s}{2}},q^{\frac{t}{2}+v+j},q^{\frac{t}{2}-s-v-j}\right).$
@@
$\pi_{q^{t},q^{s}}^{(N)}\left(y_{j}^{(N),v}(t),y_{0}^{(N),v}(s)\right)$ $=\frac{(q^{-2v-s},-q^{w+t+j+v},-q^{w-j-v},q^{t-s})_{\infty}}{(-q^{w-v},q^{t-s+j},q^{-2v-s-j},-q^{v+w+t})_{\infty}},$
$\pi_{q^{t},q^{s}}^{(N)}\left(y_{j}^{(N),v},y_{k}^{(N),v}\right)$ $=\pi_{q^{t},q^{s}}^{(N)}\left(y_{j}^{(N),v},y_{0}^{(N),v}\right)\times M_{k},$
$M_{k}$ $=\frac{(q^{2v+s},-q^{v+w+s},q^{2v+t+j},q^{-j})_{k}(1-q^{2v+2k+s})}{(q,-q^{1+v-w},q^{1+s-t-j},q^{1+2v+s+j})_{k}(1-q^{2v+s})(-q^{v+w+t-1})^{k}}.$
@@
$\mathcal{A}_{N}(j)$ $:=\mathcal{A}^{+}_{N}\left(-2v-s,t-s\right)+\mathcal{A}^{-}\left(w+t+j+v,w-j-v\right)$
$-\mathcal{A}^{+}\left(t-s+j,-2v-s-j\right)-\mathcal{A}^{-}\left(w-v,v+w+t\right)$
@@
$\pi_{q^{t},q^{s}}^{(N),d,d}\left(y_{j}^{(N),v}(t),y_{0}^{(N),v}(s)\right)=\frac{\Gamma(t-s+j)\Gamma(-2v-s-j)}{\Gamma(-2v-s)\Gamma(t-s)}e^{\textsc{Error}_{t,s}^{(N),d,d}(y_{j}^{(N),v}(t),y_{0}^{(N),v}(s))},$
@@
$M_{k}=\frac{[2v+s,2v+t+j,-j]_{k}(2v+2k+s)}{[1,1+s-t-j,1+2v+s+j]_{k}(2v+s)}(-1)^{k}e^{\overline{\textsc{Error}}_{t,s}^{(N),d,d}(y_{j}^{(N),v}(t),y_{k}^{(N),v}(s))}.$
@@
$|\overline{\textsc{Error}}_{t,s}^{(N),d,d}(y_{j}^{(N),v}(t),y_{k}^{(N),v}(s))|\leq CN^{-\chi}.$
We can also deal with the limit of the ration of $\widehat{\pi}_{t}^{(N),d}(y)$ and $\widehat{\pi}_{s}^{(N),d}(x)$ using our marginal computations.
$\lim_{N\to\infty}\frac{\widehat{\pi}_{t}^{(N),d}(\widehat{y}_{j}^{(N),v})}{\widehat{\pi}_{s}^{(N),d}(\widehat{y}_{j}^{(N),v})}$ $=\lim_{N\to\infty}\frac{N^{u+v}\widehat{\pi}_{t}^{(N),d}(\widehat{y}_{j}^{(N),v}(t))}{N^{u+v}\widehat{\pi}_{s}^{(N),d}(\widehat{y}_{k}^{(N),v}(s))}=\frac{\mathfrak{p}_{t}^{d}(x_{j}^{v}(t))}{\mathfrak{p}_{s}^{d}(x_{j}^{v}(s))}.$
@@
$\lim_{N\to\infty}\widehat{\pi}_{s,t}^{(N),d,d}(\widehat{y}_{j}^{(N),v},\widehat{y}_{k}^{(N),v})=\mathfrak{p}_{s,t}(x_{k}^{v}(s),x_{j}^{v}(t)).$
@@
$\widehat{\pi}_{s,t}^{(N),d,d}\left(\widehat{y}_{j}^{(N),u}(s),\widehat{y}_{k}^{(N),v}(t)\right)$ $=\pi_{q^{t},q^{s}}^{(N),d,d}\left(y_{k}^{(N),v}(t),y_{j}^{(N),u}(s)\right)\frac{\pi_{q^{t}}^{(N),d}\left(y_{k}^{(N),v}(t)\right)}{\pi_{q^{s}}^{(N),d}\left(y_{j}^{(N),u}(s)\right)},$
$\pi_{q^{t},q^{s}}^{(N),d,c}\left(y_{k}^{(N),v}(t),V\right)$ $=AW\left(V;q^{v+\frac{s}{2}},-q^{w+\frac{s}{2}},q^{\frac{t-s}{2}+\left(k+v+\frac{t}{2}\right)},q^{\frac{t-s}{2}-\left(k+v+\frac{t}{2}\right)}\right),$